on his final exam. gareth got 42 out of 45 questions correct. what fraction represents correct answers? what fraction represents incorrect answers?
Answer by richard1234(7193)

He got 42 correct and 3 wrong. So 42/45 = 14/15 represents the correct answers and 3/45 = 1/15 represents the incorrect answers.